OpenStudy (anonymous):

How do I graph a circle 2x^2+2y^2=8?

jimthompson5910 (jim_thompson5910):

2x^2+2y^2=8 2(x^2+y^2)=8 x^2+y^2 = 8/2 x^2+y^2 = 4

jimthompson5910 (jim_thompson5910):

Does that help?

OpenStudy (anonymous):

@jim_thompson5910 Yes, by doing that, I can see and understand the process of solving it.

mathslover (mathslover):

Sorry :(

OpenStudy (anonymous):

Oh that is true, so the radius is 2 right? @mathslover

jimthompson5910 (jim_thompson5910):

its fine

jimthompson5910 (jim_thompson5910):

yes sakigirl

OpenStudy (anonymous):

Oh okay, thank you, guys :)